OpenStudy (anonymous):

PLEASE HELP Use the Binomial Theorem to find the binomial expansion of the expression. (x – 3)^5

OpenStudy (badhi):

Do you know the binomial theorem?

OpenStudy (anonymous):

no not to clear on it

OpenStudy (badhi):

It's a theorem that enables you to expand expressions like $$(a+b)^n$$ normally we know how to expand these kind of expressions when \[n=2,3\], but for larger n values we can use this theorem. \[(a+b)^n= \sum \limits_{r=0}^{n}\binom{n}{r}a^rb^{n-r}\] where, \[\binom{n}{r}=\frac{n!}{r!(n-r)!}\] so, just substitute values a,b and n to get your answer. for more information, http://en.wikipedia.org/wiki/Binomial_theorem